# Spin-operator form factors of the critical Ising chain and their finite volume scaling limits

### v2

@@ 2.1 Inversion and summation
$W_{\beta z}=\frac{1}{\sqrt{\beta}-\sqrt{z}}+|0\rangle_{\beta}\langle v|_{z}\ ,$ (129)
$W^{\prime}_{\beta z}=\frac{1}{-\sqrt{\beta}-\sqrt{z}}-|0\rangle_{\beta}\langle v|_{v}\ ,$ (130)
@@
$W=T+|0\rangle\langle v|\ ,\ W^{\prime}=H-|0\rangle\langle v|\ .$ (133)
@@
$\langle v|T^{-1}|0\rangle=\frac{i-1}{2}\left(1-(HT^{-1})_{11}\right)\ .$ (137)
@@
$W^{-1}=T^{-1}-\frac{1}{1+\langle 0|T^{-1}|v\rangle}T^{-1}|0\rangle\langle v|T^{-1}\ .$ (141)
@@
which follows from the representation Eq. (75) and the fact that $\hat{\Gamma}_{N}(\beta)=\hat{\Gamma}(\bar{\beta})$ when $\beta\in S_{+}$. Using this relation, we obtain the desired overlapping matrix in Eq. (109)
@@ 2.2 The Wick contraction
${\cal M}_{2k+1,2l}(\beta_{i},z_{j})={\cal M}_{2k+1,2l}=(-1)^{l}\langle\prod_{i=1}^{2k+1}c_{\beta_{i}}\hat{\sigma}\prod_{j=1}^{2l}c_{z_{j}}^{\dagger}\rangle\ .$ (165)
The extra factor of $(-1)^{l}$ is equivalent to the redefinition
$c_{z}^{\dagger}\rightarrow-ic_{z}^{\dagger}\ .$ (166)
@@
$(-1)^{l}\langle\prod_{i=1}^{2k+1}c_{\beta_{i}}\hat{\sigma}\prod_{j=1}^{2l}c_{z_{j}}^{\dagger}\rangle={\cal M}_{2k+1,2l}(\beta_{i},z_{j})$
@@ 2.3 The vacuum overlap
$z_{0}=e^{i\frac{\pi}{N}},z_{1}=e^{-i\frac{\pi}{N}}....\ ,z_{2k}=e^{i\frac{2k+1}{N}},z_{2k+2}=e^{-i\frac{2k+1}{N}}\ ,\ 0\leq k\leq\frac{N-2}{2}\ .$ (182)
@@
$\det T^{2}=iN^{2N}P^{-1}\ ,\ \det T=\pm\frac{1}{\sqrt{2P}}(1+i)N^{N}$ (197)
Thus one has
@@
We should also mention that instead of using the Binet’s integral Eq. (96), if one perform the sum in the Barnes representation Eq. (97), one can obtain the following Barnes representation for the Glashier constant
@@
To obtain this form, one might need to use the the functional equation for $\zeta(u)$ to change its argument from $1+u$ to u-u. Notice that the integrand is symmetric under $u\rightarrow-u$.
@@
$-\frac{\pi^{2}}{48N^{2}}\sum_{n=0}^{\infty}\left(\left(n+\frac{1}{2}\right)^{-u}-\left(n+1\right)^{-u}\right)\bigg|_{u\rightarrow-1}=-\frac{\pi^{2}}{384N^{2}}\ ,$ (214)
$\frac{7\pi^{4}}{23040N^{4}}\bigg(\sum_{m\in Z_{\geq 0}+\frac{1}{2}}\frac{1-8m^{2}}{m^{u}}-\sum_{n\in Z_{\geq 1}}\frac{1-8n^{2}}{n^{u}}\bigg)\bigg|_{u\rightarrow 1}=\frac{7\pi^{4}}{92160N^{4}}\ .$ (215)
@@ 2.4 Scaling limit of the two-point correlator
This is exactly the large distance asymptotics of the infinite-volume critical chain. As such, the simple fixed-point scaling scenario defined in McCoy:2000gw continues to hold in this example, if one believes that $\det(1+{\cal K})=(1-w)^{-\frac{1}{8}}$.
@@
This is again consistent with the standard CFT formula DiFrancesco:1997nk.
@@ 3 Comparison with literature and further comments
First, in the introduction, we mentioned that the massless limits of the finite volume massive form factors in Fonseca:2001dc were rarely mentioned, so the first thing we would like to comment are the leg functions in that reference. In the notation of Fonseca:2001dc, the function $\kappa(\theta)$ reads
@@
The first line is essentially the integral given in Gabai:2019ryw, while in the second line, we have averaged with the $k\rightarrow-k$ integrand and changed to $E^{\prime}=\sqrt{k^{2}+\mu^{2}}$. In this form, it finally becomes transparent that there is a non-trivial $\mu\rightarrow 0$ limit at fixed ELEL
@@
This is exactly the Binet’s integral Eq. (96). As such, at the quantization levels $EL=2\pi|n|$ of the R sector, and at $EL=2\pi|m|$ ($m\in Z+\frac{1}{2}$) for the NS sector, the leg functions in Fonseca:2001dc evaluate to
@@
$\frac{e^{-\kappa(\theta_{m})}}{\sqrt{\mu L\cosh\theta_{m}}}\rightarrow\frac{\sqrt{2\pi}\Gamma(m+1)}{2\pi m\Gamma(m+\frac{1}{2})}=\tau\left(m-\frac{1}{2}\right)$ (237)
@@
again consistent with the $d(1)=1$. Two particle functions clearly simplify to the $\Pi$ factors in Eq. (104) and Eq. (105). Moreover, in Appendix. A, we show that normalization function $\bar{\sigma}S(mL)$ of Fonseca:2001dc has the following small $m$ expansion
$\bar{\sigma}S(mL)=\left(\frac{2\pi}{L}\right)^{\frac{1}{8}}\exp\bigg(\frac{\ln 2}{2\pi}mL-\frac{\ln^{2}2}{4\pi^{2}}(mL)^{2}-\frac{\zeta_{3}}{16\pi^{3}}(mL)^{3}+{\cal O}((mL)^{4})\bigg)\ .$ (240)
Given this, we conclude that the massless limit of the massive matrix elements in Fonseca:2001dc, up to overall phase factors, exactly reduce to $\left(\frac{2\pi}{L}\right)^{\frac{1}{8}}$ multiplying our Eq. (104) and Eq. (105) for $\hat{\sigma}$. The normalization of Fonseca:2001dc ensures the small-$r$ limit of the massless two point correlator to be $r^{-\frac{1}{4}}$ without extra factors.
Second, in the introduction, we have mentioned that in Gehlen_2008, the authors have omitted the $g=1$ point in their discussions, but this does not rule out the possibility that the results therein also applies at the $g=1$ by taking limit. Let’s show that this is indeed the case. The major quantity to check are the products of their leg function. In our notation, the leg function in Gehlen_2008 reads
@@
where we have again used $p(z)=e^{-\frac{i\pi}{4}}|p(z)|$, Eq. (61) and Eq. (78). As such, we have found that all the leg functions of the $g\rightarrow 1^{+}$ limit in Gehlen_2008 agree with our results. For the two-particle functions, it is also easy to check that they agree with our expressions. Finally, for the overall normalization factor $\xi\xi_{T}$, notice that the $\xi=(g^{2}-1)^{\frac{1}{4}}$ again cancels with the $(2\epsilon(1))^{\frac{1}{4}}$ in the denominator, and the $g\rightarrow 1^{+}$ limit is again finite. Moreover, using the relations above, one can compute
@@
where we have used the relation p(1)q(1)=2ip(1)q(1)=-2i in Eq. (71). As such, all the form factor squared in Gehlen_2008; Iorgov:2010yv, in the proper $g\rightarrow 1^{+}$ limit approached from the disordered side, agree with the results derived directly at the critical point.
@@
we have not yet verified this analytically. However, we have noticed that the form factor expansion of this determinant is similar to the infinite-volume form factor expansion Lyberg_2007 of the Ising chain or diagonal Ising correlators away from the critical point, and at zero distance. For example, the ${\rm Tr}{\cal K}$ equals to
@@
where $K(w)$ is the complete Elliptic integral of the second Kind. The $\frac{K(w)}{\pi}$ is exactly half of the one particle form factor contribution $f^{(1)}_{0,0}(t)$ to the diagonal correlator at zero separation Boukraa:2006bt. If such relations could be established at high orders, then the identity Eq. (249) becomes understandable: $\sigma_{00}^{2}=1$, while the form factor expansion in Boukraa:2006bt contains an overall $(1-t)^{\frac{1}{4}}$. This implies that all the form factors at zero separation must sum to $(1-t)^{-\frac{1}{4}}$. It is also possible that differential equations in $w$ could be established, similar to the Painleve VI for the parameter derivatives 10.3792/pjaa.56.405. Finally, it is also interesting to see how the form factors can be summed to the more general toroidal correlators at a finite temperature.
@@ Acknowledgements.
The author thanks Zoltán Bajnok for constructive discussions and introduction to the references Yurov:1991my; Fonseca:2001dc.
@@ Appendix A The $s\rightarrow 0^{+}$ expansion of the crossover function
In this appendix, we compute analytically the small $s\rightarrow 0^{+}$ expansion of the function $g_{I}(s)$ of Sachdev:1995bf. We start from the definition with $s>0$
@@
In particular, the leading term exactly reproduces the constant ${\cal C}_{1}$ in Sachdev:1995bf extracted numerically. At higher orders, there will be $\zeta_{2k+1}\zeta_{2l+1}$ type terms. The expansion has a convergence radius equals to $\pi$.
Given the expansion of $\ln g_{I}(s)$, the expansion of the function $S(mL)$ in Fonseca:2001dc can also be found using the relation
$\ln S(mL)=\frac{1}{2}\ln g_{I}\left(mL\right)-\frac{1}{8}\ln mL\ ,$ (262)
which is equivalent to Eq. (A13) of Fonseca:2001dc. Especially, using the expansion Eq. (261), in the massless limit the $\bar{\sigma}m^{\frac{1}{8}}S(mL)$ is finite
$\lim_{m\rightarrow 0}m^{\frac{1}{8}}2^{\frac{1}{12}}e^{-\frac{1}{8}}A^{\frac{3}{2}}S(mL)=\frac{2^{\frac{1}{12}}e^{-\frac{1}{8}}A^{\frac{3}{2}}}{L^{\frac{1}{8}}}(\pi e)^{\frac{1}{8}}2^{\frac{1}{24}}A^{-\frac{3}{2}}=\left(\frac{2\pi}{L}\right)^{\frac{1}{8}}\ .$ (263)
Using the Eq. (2.15) of Fonseca:2001dc and the massless limits of their leg functions discussed in Sec. 3, this implies that the overall normalization of the spin operator in Fonseca:2001dc is exactly $\left(\frac{2\pi}{L}\right)^{\frac{1}{8}}$ times our $\hat{\sigma}$. In particular, at small $r$, the two point correlator approaches a simple power-law $r^{-\frac{1}{4}}$ without extra factors.
@@
The integral here is similar to that of the Ising connecting computation Tracy:1990tn. Using $\psi(x)=(\ln\Gamma(x))^{\prime}$ and partial integrating as in Tracy:1990tn, one has

### v3

@@ 2.1 Inversion and summation
$W_{\beta z}=\frac{\sqrt{\beta}+\sqrt{z}}{\beta-z}+\delta_{\beta 1}\frac{i-1}{1-z}\equiv\frac{1}{\sqrt{\beta}-\sqrt{z}}+|0\rangle_{\beta}\langle v|_{z}\ ,$ (132)
$W^{\prime}_{\beta z}=-\frac{\sqrt{\beta}-\sqrt{z}}{\beta-z}-\delta_{\beta 1}\frac{i-1}{1-z}\equiv\frac{1}{-\sqrt{\beta}-\sqrt{z}}-|0\rangle_{\beta}\langle v|_{z}\ ,$ (133)
@@
$W=T+|0\rangle\langle v|\ ,\ W^{\prime}=H-|0\rangle\langle v|\ ,v_{z}=\frac{i-1}{2}\left(T_{1z}-H_{1z}\right)\ .$ (136)
@@
$\langle v|T^{-1}|0\rangle=v_{z}(T^{-1})_{z1}=\frac{i-1}{2}\left(1-(HT^{-1})_{11}\right)\ .$ (140)
@@
$W^{-1}=T^{-1}-\frac{1}{1+\langle v|T^{-1}|0\rangle}T^{-1}|0\rangle\langle v|T^{-1}\ .$ (144)
@@
which follows from the representation Eq. (75) and the fact that $\hat{\Gamma}_{N}({\cal Z})=\hat{\Gamma}(\bar{\cal Z})$ when ${\cal Z}\in S_{\pm}$. Using this relation, we obtain the desired overlapping matrix in Eq. (112)
@@ 2.2 The Wick contraction
${\cal M}_{2k+1,2l}(\{\beta_{i}\},\{z_{j}\})={\cal M}_{2k+1,2l}=(-1)^{l}\langle\prod_{i=1}^{2k+1}c_{\beta_{i}}\hat{\sigma}\prod_{j=1}^{2l}c_{z_{j}}^{\dagger}\rangle\ .$ (168)
The extra factor of $(-1)^{l}$ can be achieved through the redefinition
$c_{z}^{\dagger}\rightarrow-ic_{z}^{\dagger}\ ,\ c_{z}\rightarrow ic_{z}\ ,$ (169)
@@
$(-1)^{l}\langle\prod_{i=1}^{2k+1}c_{\beta_{i}}\hat{\sigma}\prod_{j=1}^{2l}c_{z_{j}}^{\dagger}\rangle={\cal M}_{2k+1,2l}(\{\beta_{i}\},\{z_{j}\})$
@@ 2.3 The vacuum overlap
$z_{0}=e^{i\frac{\pi}{N}},z_{1}=e^{-i\frac{\pi}{N}}....\ ,z_{2k}=e^{i\frac{2k+1}{N}\pi},z_{2k+2}=e^{-i\frac{2k+1}{N}\pi}\ ,\ 0\leq k\leq\frac{N-2}{2}\ .$ (185)
@@
$\det T^{2}=iN^{2N}P^{-1}\ ,\ \det T=\pm\frac{1}{\sqrt{2P}}(1+i)N^{N}\ ,$ (200)
leading to
@@
We should also mention that instead of using the Binet’s integral Eq. (96), if one perform the sum in the Barnes representation Eq. (97), one can obtain the following Barnes representation for the Glaisher constant
@@
Notice that the integrand is symmetric under $z\rightarrow-z$. To be more precise, one first write
$\sum_{i=0}^{\infty}\bigg(\mathfrak{b}(\pi(2i+1))-\mathfrak{b}(2\pi(i+1))\bigg)$
$=\sum_{i=0}^{\infty}\bigg(\ln\sqrt{i+\frac{1}{2}}+\ln\frac{\Gamma\left(i+1\right)}{\Gamma(i+\frac{3}{2})}-\ln\sqrt{i+1}-\ln\frac{\Gamma\left(i+\frac{3}{2}\right)}{\Gamma(i+2)}\bigg)$
$=\ln\left(\pi^{\frac{1}{4}}G\left(\frac{1}{2}\right)G\left(\frac{3}{2}\right)\right)\ .$ (216)
where the $\mathfrak{b}(\alpha)$ is defined in Eq. (96), and we have used Eq. (212). Then, due to the absolute convergence, we can sum under the Barnes integral Eq. (97). This leads to
$\sum_{i=0}^{\infty}\bigg(\mathfrak{b}(\pi(2i+1))-\mathfrak{b}(\pi(2i+2))\bigg)$
$=-\int_{0<{\rm Re}(u)<1}\frac{du}{2\pi i}\left(1-\frac{1}{2^{u+1}}\right)\frac{\Gamma(u)\zeta(u+1)}{\cos\frac{\pi u}{2}}(2\pi)^{-u}\times(2^{u}-2)\zeta(u)\ .$ (217)
Finally, we use the functional equation of the $\zeta$ function to change $\zeta(u+1)$ to $\zeta(-u)$ and obtain Eq. (215). In Appendix. A, we will use this Barnes integral to find the small mass expansion of a related crossover function.
@@
$-\frac{\pi^{2}}{48N^{2}}\left(\sum_{m\in Z_{\geq 0}+\frac{1}{2}}\frac{1}{m^{u}}-\sum_{n\in Z_{\geq 1}}\frac{1}{n^{u}}\right)\bigg|_{u\rightarrow-1}=-\frac{\pi^{2}}{384N^{2}}\ ,$ (219)
$\frac{7\pi^{4}}{23040N^{4}}\bigg(\sum_{m\in Z_{\geq 0}+\frac{1}{2}}\frac{1-8m^{2}}{m^{u}}-\sum_{n\in Z_{\geq 1}}\frac{1-8n^{2}}{n^{u}}\bigg)\bigg|_{u\rightarrow-1}=\frac{7\pi^{4}}{92160N^{4}}\ .$ (220)
@@ 2.4 Scaling limit of the two-point correlator
This is exactly the large distance asymptotics of the infinite-volume critical chain. As such, the simple fixed-point scaling scenario defined in McCoy (2000) continues to hold in this example, if one believes that $\det(1+{\cal K})=(1-w)^{-\frac{1}{8}}$.
@@
This is again consistent with the standard CFT formula Di Francesco et al. (1997).
@@ 3 Comparison with literature and further comments
First, in the introduction, we mentioned that the massless limits of the finite volume massive form factors in Fonseca and Zamolodchikov (2001) were rarely mentioned, so the first thing we would like to comment are the leg functions in that reference. In the notation of Fonseca and Zamolodchikov (2001), the function $\kappa(\theta)$ reads
@@
The first line is essentially the integral given in Gabai and Yin (2022), while in the second line, we have averaged with the $k\rightarrow-k$ integrand and changed to $E^{\prime}=\sqrt{k^{2}+\mu^{2}}$. In this form, it finally becomes transparent that there is a non-trivial $\mu\rightarrow 0$ limit at fixed ELEL
@@
This is exactly the Binet’s integral Eq. (96). As such, at the quantization levels $EL=2\pi|n|$ of the R sector, and at $EL=2\pi|m|$ ($m\in Z+\frac{1}{2}$) for the NS sector, the leg functions in Fonseca and Zamolodchikov (2001) evaluate to
@@
$\frac{e^{-\kappa(\theta_{m})}}{\sqrt{\mu L\cosh\theta_{m}}}\rightarrow\frac{\sqrt{2\pi}\Gamma(m+1)}{2\pi m\Gamma(m+\frac{1}{2})}=\tau\left(m-\frac{1}{2}\right)\ ,$ (242)
@@
again consistent with the $d(1)=1$. Two particle functions clearly simplify to the $\Pi$ factors in Eq. (107) and Eq. (108). Moreover, in Appendix. A, we show that normalization function $\bar{\sigma}S(\mu L)$ of Fonseca and Zamolodchikov (2001) has the following small $\mu$ expansion
$\bar{\sigma}S(\mu L)=\left(\frac{2\pi}{L}\right)^{\frac{1}{8}}\exp\bigg(\frac{\ln 2}{2\pi}\mu L-\frac{\ln^{2}2}{4\pi^{2}}(\mu L)^{2}-\frac{\zeta_{3}}{16\pi^{3}}(\mu L)^{3}+{\cal O}((\mu L)^{4})\bigg)\ .$ (245)
Given this, we conclude that the massless limit of the massive matrix elements in Fonseca and Zamolodchikov (2001), up to overall phase factors, exactly reduce to $\left(\frac{2\pi}{L}\right)^{\frac{1}{8}}$ multiplying our Eq. (107) and Eq. (108) for $\hat{\sigma}$. The normalization of Fonseca and Zamolodchikov (2001) ensures the small-$r$ limit of the massless two point correlator to be $r^{-\frac{1}{4}}$ without extra factors.
Second, in the introduction, we have mentioned that in Gehlen et al. (2008), the authors have omitted the $g=1$ point in their discussions, but this does not rule out the possibility that the results therein also applies at the $g=1$ by taking limit. Let’s show that this is indeed the case. The major quantity to check are the products of their leg function. In our notation, the leg function in Gehlen et al. (2008) reads
@@
where we have again used $p(z)=e^{-\frac{i\pi}{4}}|p(z)|$, Eq. (61) and Eq. (78). As such, we have found that all the leg functions of the $g\rightarrow 1^{+}$ limit in Gehlen et al. (2008) agree with our results. For the two-particle functions, it is also easy to check that they agree with our expressions. Finally, for the overall normalization factor $\xi\xi_{T}$, notice that the $\xi=(g^{2}-1)^{\frac{1}{4}}$ again cancels with the $(2\epsilon(1))^{\frac{1}{4}}$ in the denominator, and the $g\rightarrow 1^{+}$ limit is again finite. Moreover, using the relations above, one can compute
@@
where we have used the relation p(1)q(1)=2ip(1)q(1)=-2i in Eq. (71). As such, all the form factor squared in Gehlen et al. (2008); Iorgov et al. (2011), in the proper $g\rightarrow 1^{+}$ limit approached from the disordered side, agree with the results derived directly at the critical point.
@@
we have not yet verified this analytically. However, we have noticed that the form factor expansion of this determinant is similar to the infinite-volume form factor expansion Lyberg and McCoy (2007) of the Ising chain or diagonal Ising correlators away from the critical point, and at zero distance. For example, the ${\rm Tr}{\cal K}$ equals to
@@
where $K(w)$ is the complete Elliptic integral of the second Kind. The $\frac{K(w)}{\pi}$ is exactly half of the one particle form factor contribution $f^{(1)}_{0,0}(t)$ to the diagonal correlator at zero separation Boukraa et al. (2007). If such relations could be established at high orders, then the identity Eq. (254) becomes understandable: $\sigma_{00}^{2}=1$, while the form factor expansion in Boukraa et al. (2007) contains an overall $(1-t)^{\frac{1}{4}}$. This implies that all the form factors at zero separation must sum to $(1-t)^{-\frac{1}{4}}$. It is also possible that differential equations in $w$ could be established, similar to the Painleve VI for the parameter derivatives Jimbo and Miwa (1980). Finally, it is also interesting to see how the form factors can be summed to the more general toroidal correlators at a finite temperature.
@@ Acknowledgements.
The author thanks Zoltán Bajnok for constructive discussions and introduction to the references Yurov and Zamolodchikov (1991); Fonseca and Zamolodchikov (2001).
@@ Appendix A The $s\rightarrow 0^{+}$ expansion of the crossover function
In this appendix, we compute analytically the small $s\rightarrow 0^{+}$ expansion of the function $g_{I}(s)$ of Sachdev (1996). We start from the definition with $s>0$
@@
In particular, the leading term exactly reproduces the constant ${\cal C}_{1}$ in Sachdev (1996) extracted numerically. At higher orders, there will be $\zeta_{2k+1}\zeta_{2l+1}$ type terms. The expansion has a convergence radius equals to $\pi$.
Given the expansion of $\ln g_{I}(s)$, the expansion of the function $S(\mu L)$ in Fonseca and Zamolodchikov (2001) can also be found using the relation
$\ln S(\mu L)=\frac{1}{2}\ln g_{I}\left(\mu L\right)-\frac{1}{8}\ln\mu L\ ,$ (267)
which is equivalent to Eq. (A13) of Fonseca and Zamolodchikov (2001). Especially, using the expansion Eq. (266), in the massless limit the $\bar{\sigma}\mu^{\frac{1}{8}}S(\mu L)$ is finite
$\lim_{\mu\rightarrow 0}\mu^{\frac{1}{8}}2^{\frac{1}{12}}e^{-\frac{1}{8}}A^{\frac{3}{2}}S(\mu L)=\frac{2^{\frac{1}{12}}e^{-\frac{1}{8}}A^{\frac{3}{2}}}{L^{\frac{1}{8}}}(\pi e)^{\frac{1}{8}}2^{\frac{1}{24}}A^{-\frac{3}{2}}=\left(\frac{2\pi}{L}\right)^{\frac{1}{8}}\ .$ (268)
Using the Eq. (2.15) of Fonseca and Zamolodchikov (2001) and the massless limits of their leg functions discussed in Sec. 3, this implies that the overall normalization of the spin operator in Fonseca and Zamolodchikov (2001) is exactly $\left(\frac{2\pi}{L}\right)^{\frac{1}{8}}$ times our $\hat{\sigma}$. In particular, at small $r$, the two point correlator approaches a simple power-law $r^{-\frac{1}{4}}$ without extra factors.
@@
The integral here is similar to that of the Ising connecting computation Tracy (1991). Using $\psi(x)=(\ln\Gamma(x))^{\prime}$ and partial integrating as in Tracy (1991), one has